OpenStudy (anonymous):

Factor P completely. x^4-9x^2-10

OpenStudy (anonymous):

(x^2+1)(x^2-10); x=sqrt(10), -sqrt(10)

OpenStudy (majikdusty):

If you look at the equation, it is quadratic like. if you set x^2 = u, you turn it into the equation u^2-9u-10 which factors to (u-10) and (u+1) then you can plug x^2 back in for u this gives you (x^2-10)(x^2+1) (x^2-10) is a difference of squares (if you are familiar with this term). This factors to (x+sqrt(10))(x-sqrt(10)) There are also imaginary roots for this problem